OpenStudy (moonlitfate):

Find the derivative of the function. y = (x^2-4)^(1/2) - 2 arcsec(x/2)

OpenStudy (fibonaccichick666):

so, any initial thoughts?

OpenStudy (moonlitfate):

Well, I was able to find the derivative of arcsec(x/2) -- without simplifying.

OpenStudy (moonlitfate):

@FibonacciChick666 -- this what I got for that part.

OpenStudy (fibonaccichick666):

ok, and what did you get?

OpenStudy (moonlitfate):

\[\frac{ d }{ dx }[arsec \frac{ x }{ 2 }] = \frac{ \frac{ 1 }{ 2 } }{ |\frac{ x }{ 2 }|\sqrt{(\frac{ x }{ 2 })^2-1}}\]

OpenStudy (moonlitfate):

*arcsec

OpenStudy (fibonaccichick666):

ok i can buy that, now how about the first part?

OpenStudy (moonlitfate):

I know it's the chain rule (it still confuses me a bit), but I would have to re-write it. \[y = (x^2-4)^{1/2}\]

OpenStudy (moonlitfate):

\[y \prime = \frac{ 1 }{ 2 } * (x^2-4)^{-1/2} * 2x\] I think maybe? :/ Am I right @FibonacciChick666

OpenStudy (fibonaccichick666):

yup yup

OpenStudy (fibonaccichick666):

it's that simple

OpenStudy (fibonaccichick666):

now, you can take a longer approach to make it easier on your brain though if you say let \[u=x^2-4\]\[du=2x\] so now you have \[y=u^\frac{1}{2}\]

OpenStudy (fibonaccichick666):

if that helps

OpenStudy (moonlitfate):

Hah, yeah. It's trying to get that final answer. So many algebraic things that are easy to mess up.

OpenStudy (fibonaccichick666):

true true

OpenStudy (fibonaccichick666):

but you have it, just be confident

OpenStudy (fibonaccichick666):

if you're worrid about messing up, always do the substitution on the side

OpenStudy (moonlitfate):

Yeah, that's what I've been doing. It's tedious and time-consuming, but I've found it helps.

OpenStudy (fibonaccichick666):

it does, and it will help when you get to integration by parts to be in the habit

OpenStudy (fibonaccichick666):

you'll get better with practice :)

OpenStudy (fibonaccichick666):

medal?